From a sample of 65 meth addicts, you find that the mean number of tattoos is 17.3, with a sample standard deviation of 6.8. Test the null hypothesis that your sample came from a population whose mean number of tattoos is 15 against the alternative hypothesis that Mu is less than 15. Use alpha = .01.

Formally state the null and the alternative hypothesis for this test (3 pts)

What is the critical value(s) for this test? (3 pts)

Interpret your decision regarding the null hypothesis. (3 pts)

If necessary, compute 99% confidence intervals for the population mean and interpret. (3 pts)
Answer #1

solution- Given that, x = 17.3 G= 6-8 n= 65 l= 15 L- o-ol df=n-1=65-1=64 Hoiu=15 Ha! ll <15 this is left-tailed test Test stak X = 17.3 s = 6.8 .0 n= 65 die n-1365-1 =64 At aq olo confidence level the tis L=1-49 olo dol-0.99-0-01 V =001. - 01005 +212